OpenStudy (lexy148):

what times what equals 60 and adds up to 15?

OpenStudy (anwaarullah):

12*5=60 10*6=60 is it??

OpenStudy (anonymous):

We can find out "what times what" equals 60 by finding the square root. \[\sqrt{60}\] = 7.745 Then we can multiply 7.745 by 2 which ≈ 15.

OpenStudy (lexy148):

thx